OpenStudy (anonymous):

What is the greatest common factor of 20x4 and 6x3? A. x3 B. x4 C. 2x3 D. 2x4

OpenStudy (mathstudent55):

Look at the numbers first: 20 and 6. What is their GCF?

OpenStudy (anonymous):

is it 2 ?

OpenStudy (mathstudent55):

20 = 2 * 2 * 5 6 = 2 * 3 The only factor they have in common is 2.

OpenStudy (anonymous):

at least i was right , what's next ?

OpenStudy (mathstudent55):

Correct. Now look at the variable parts: x^4 and x^3. x^4 = x * x * x * x x^3 = x * x * x How many x's do they have in common?

OpenStudy (anonymous):

if were counting the x in front of the number's then they have 4 in common but if not they only have 3 in common .

OpenStudy (mathstudent55):

No. It's true that the first one has 4 x's and the second one has 3 x's. How many they have in common means the _highest_ number of x's they both have. They both have 3 x's, but they both don't have 4, so they have 3 x's in common. That means they have x^3 in common.

OpenStudy (anonymous):

so the answer is x^3

OpenStudy (mathstudent55):

Yes, for the variables in common Now you put together both parts (the number in common and the variables in common)to form your answer: The two given terms have 2 in common and x^3 in common. The GCF is 2x^3